### v3

# EEGDM: Learning EEG Representation with Latent Diffusion Model

Abstract.
Recent advances in self-supervised learning for EEG representation have largely relied on masked reconstruction, where models are trained to recover randomly masked signal segments. While effective at modeling local dependencies, such objectives are inherently limited in capturing the global dynamics and long-range dependencies essential for characterizing neural activity. To address this limitation, we propose EEGDM, a novel self-supervised framework that leverages latent diffusion models to generate EEG signals as an objective. Unlike masked reconstruction, diffusion-based generation progressively denoises signals from noise to realism, compelling the model to capture holistic temporal patterns and cross-channel relationships. Specifically, EEGDM incorporates an EEG encoder that distills raw signals and their channel augmentations into a compact representation, acting as conditional information to guide the diffusion model for generating EEG signals. This design endows EEGDM with a compact latent space, which not only offers ample control over the generative process but also can be leveraged for downstream tasks. Experimental results show that EEGDM (1) reconstructs high-quality EEG signals, (2) learns robust representations, and (3) achieves competitive performance across diverse downstream tasks, thus exploring a new direction for self-supervised EEG representation learning.
Electroencephalography, Representation Learning, Self-supervised Learning, Diffusion Transformers
conference: ; ; copyright: noneisbn: 978-1-4503-XXXX-X/2018/06submissionid: 5469ccs: Computing methodologies Machine learningccs: Human-centered computing Human computer interaction (HCI)$c$s: Information systems Multimedia information systems
1. Introduction
Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes. The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies (Mohsenvand et al., 2020). Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components (Zander and Kothe, 2011).
Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data (Jia et al., 2020, 2021; Jiang et al., 2024a). Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models. Task-specific models face significant challenges in generalization due to substantial variations across datasets in channel configurations, sampling rates, and EEG recording time lengths. These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability. For multiple tasks, three large EEG models, EEGPT (Wang et al., 2024), LaBraM (Jiang et al., 2024c), and CBraMod (Wang et al., 2025) have recently been proposed. These models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks. However, masked reconstruction has limited capacity in capturing the global dynamics and long-term dependencies essential for characterizing neural activity.
In this paper, we introduce the Diffusion Model (DM) (Ho et al., 2020; Dhariwal and Nichol, 2021) into EEG representation learning and propose a novel self-supervised framework called EEGDM. Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. Unlike popular masked reconstruction paradigms that primarily focus on local patches recovery, DM learns effective representations through a self-supervised denoising objective. This objective necessitates a profound understanding of the underlying data distribution, thereby compelling the latent space to preserve macro-level neural patterns. To successfully denoise a latent representation into a coherent EEG segment, the model must capture the holistic phase relationships and cross-channel correlations that define specific brain states. This makes DM potentially valuable for generic EEG representation learning. Building on this observation, our framework uses an EEG encoder that transforms input signals and their channel augmentations into compact latent representations, which condition the diffusion process. This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity (Gramfort et al., 2013; Cole and Voytek, 2019). To address these challenges, we incorporate two synergistic components into our EEGDM framework. First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information. Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones (Sanei and Chambers, 2013). This transformation mitigates the challenge of reconstructing the original noisy signal. More importantly, PCA simplifies the latent manifold, making the model more sensitive to the structural information provided by the augmented views. Meanwhile, the diverse augmented inputs prevent the model from losing critical information when training against a reduced-dimension (PCA) target. Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) (Rombach et al., 2022) and Diffusion Transformers (DiTs) (Peebles and Xie, 2023)) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space. This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks. Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks and exhibiting strong generalizability.
Our contributions are summarized as follows:
  • We propose EEGDM, a novel diffusion model-based self-supervised framework for EEG representation learning. By leveraging EEG signal generation as the self-supervised objective, EEGDM transforms the diffusion model into a powerful representation learner that captures rich EEG semantics.
  • EEGDM incorporates channel augmentation and elaborately designed PCA-based latent space operations to (1) enhance the conditional information available for DM, (2) enable robust EEG signal reconstruction, and (3) improve the cross-dataset generalizability of learned EEG representations.
  • We validate the effectiveness and generalizability of EEGDM across multiple EEG analysis tasks, demonstrating its (1) competitive performance compared to existing methods and (2) strong cross-dataset transferability and practical utility in EEG applications.
2. Related Work
2.1. Self-supervised EEG Representation Learning
Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations. This methodology has led to significant breakthroughs in natural language processing and computer vision (Kenton and Toutanova, 2019; He et al., 2022). However, its full potential in EEG analysis remains largely unexplored (Jiang et al., 2025).
Some studies (Mohsenvand et al., 2020; Banville et al., 2021) have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework (Chen et al., 2020) to time series data. Another line of work extends the mask idea to EEG analysis (Mohammadi Foumani et al., 2024). For instance, BENDR (Kostas et al., 2021) uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part. LaBraM (Jiang et al., 2024c) pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook. EEGPT (Wang et al., 2024) introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features. CBraMod (Wang et al., 2025) introduces a structurally aware EEG foundation model that separately captures heterogeneous spatial–temporal dependencies through a criss-cross transformer and achieves strong cross-dataset generalization via large-scale masked pre-training.
Unlike the above approaches, we explore self-supervised EEG representation learning via generative models. Our proposed EEGDM not only facilitates high-quality EEG generation but also enables the learning of rich and robust representations for downstream tasks.
2.2. Diffusion Model
Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics. This fundamental characteristic endows them with dual capabilities: generating high-quality samples and learning meaningful representations inherent in the data distribution. The pioneering work by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015) and subsequent advances by Ho et al. (Ho et al., 2020) established diffusion models as powerful tools, leading to widespread applications in generative tasks (Rombach et al., 2022; Peebles and Xie, 2023; Zhang et al., 2023). More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features (Mittal et al., 2023; Chen et al., 2024b; Hudson et al., 2024).
Figure 1. The overall architecture of EEGDM. The original EEG data is projected onto a latent space using PCA. Noise is added in this latent space, and the denoised latent representation is predicted by learning the conditional latent diffusion model. The prediction is then projected back to the original space by inverse PCA to obtain the generated target EEG signal. The encoder transforms the source EEG signal (including original data and augmentations) into a concise EEG representation. It is combined with the noise timestep representation and then used as conditional information to modulate the DiT block, thus guiding the denoising of the latent representation.
In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation (Huang et al., 2024; Chen et al., 2024a; Huang et al., 2025), or on a few specific supervised tasks (Kim et al., 2023, 2024). Recently, Puah et al. (Puah et al., 2025) proposed an EEG representation learning framework, also termed EEGDM. To avoid ambiguity, we refer to their method as Puah-EEGDM-SSM. Their approach integrates a Structured State-Space Model (SSM) into the diffusion pre-training paradigm to better capture the temporal dynamics of EEG signals. In contrast, our EEGDM adopts a fundamentally different architectural paradigm. Specifically, we employ a Latent Diffusion Transformer (DiT) operating on a refined latent manifold, which is specifically designed to capture global inter-channel spatial synchrony and complex phase relationships—critical neural features that are overlooked by the linear recurrence mechanisms inherent in SSM.
3. Method
Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying distributions, structures, and dynamics of the data during the generative process. Therefore, we explore self-supervised EEG representation learning via the diffusion model, and design the following framework architecture.
3.1. Framework Architecture
Our proposed self-supervised EEG representation learning framework, named EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation. Our goal is to generate high-quality EEG signals while equipping the model with a generalizable understanding of EEG data that can be effectively adapted to various downstream tasks. Notably, reconstructing raw EEG signals directly remains formidable due to their inherently low SNR. To mitigate this, we employ PCA to project raw data into a refined latent space with enhanced SNR, shifting the reconstruction target to these denoised components. By focusing on the top-K principal components, the diffusion module effectively constrains the search space to the most energetically significant neural manifolds, thereby alleviating the optimization burden. Crucially, while EEG augmentation introduces view diversity to force the extraction of invariant features, PCA provides a stable, denoised reconstruction objective. The overall architecture of EEGDM is shown in Fig. 1.
Specifically, given the original multi-channel EEG signals $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of EEG electrodes (channels) and $T$ is the total timestamp, we segment $X$ into ${\lfloor{\frac{T-t^{s}}{s^{t}}}\rfloor}+1$ samples. This segmentation assumes a timestamp length [^1] Here superscript is used; subscript is reserved for diffusion formation in subsequent sections. $t^{s}$ per sample and a stride $s^{t}$. As shown in Fig. 1, each multi-channel sample $x\in\mathbb{R}^{C\times{t^{s}}}$ is processed by PCA to obtain the latent representation $z$. Our diffusion model operates on this latent space, where noise is added and predicted. The denoised latent representation predicted by the diffusion model is projected back to the original space by inverse PCA to obtain the target EEG signal. Owing to the effectiveness of the DiT model in the latent space, we utilize it to predict the denoised latent representations. Furthermore, we introduce an EEG encoder that transforms the input source signal $x^{\prime}$ (including original EEG sample and its augmentations) into an EEG representation containing rich semantic information. The sum of this EEG representation and the noise timestep representation (i.e., timestep $t$ in the noise-adding process) encoded by the time embedding layer serves as the conditional input information for the diffusion model. This diffusion model design can utilize the semantically rich external EEG representation for the generative task, leading to a significant improvement in generative performance. Consequently, this enhances the model’s representation learning capability.
3.2. The EEG Encoder
The representation output from the EEG encoder lies at the core of EEGDM, which (1) acts as the communication bridge between the encoder and the denoiser ($\epsilon_{\theta}$), and (2) guides the denoising process. In more details, this output represents a clean source EEG signal semantically related to the target EEG data. The introduction of the encoded EEG representation is to enable hidden states of the diffusion model to predict noise-invariant and semantically meaningful EEG representations from noisy inputs, guiding the reconstruction of the target EEG data during the denoising process.
The core motivation behind the above EEG encoder design arises from the inherently under-determined nature of the denoising task. The denoiser tends to exploit any relevant knowledge or informative cues to infer and recover the original data distribution as accurately as possible. This incentivizes the encoder to distill salient commonalities between the source and target EEG signals into a compact representation. We constrain the encoded EEG representation to a low-dimensional latent space and guide the denoiser through feature modulation. This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser. Rather than reconstruction from scratch as most conventional denoisers do (Hinton and Salakhutdinov, 2006; Higgins et al., 2017), our strategy enables the encoder to learn a representation that supports latent space denoising. This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG. Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
Our EEG encoder is based on the Vision Transformer (ViT) architecture (Dosovitskiy, 2021) that operates on the sequences of tokens. We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences. We retain many of ViT’s best practices, enabling the EEG encoder to be a Transformer architecture that can process input EEG signals with various formats (different sampling rates, number of channels, and time lengths). The input of the encoder is the source EEG signal $x^{\prime}$, which consists of the original samples and the EEG channel augmentations. EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data. Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work (Mohsenvand et al., 2020) to be the most effective for extracting useful features.
Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$). The total number of patches for $x^{\prime}$ is ${\left|x^{\prime}\right|}=C{\lfloor{\frac{t^{s}}{\omega}}\rfloor}$. Then a temporal encoder (i.e., a temporal convolution block) is used to encode each EEG patch into a patch embedding $e^{\prime}=\{e^{\prime}_{j,k}\in\mathbb{R}^{d}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$, where $d$ is the dimension of the embeddings. The learnable positional embeddings are added to all input patch embeddings. Finally, the sequence of embeddings is fed directly into the Transformer network. We use average pooling on the output embeddings to obtain the EEG representation, which is then combined with the noise timestep representation to guide the denoising of the diffusion model. As illustrated in Fig. 2, we evaluate the model’s performance on various downstream tasks by appending a task-specific linear prediction head to the pre-trained encoder. For each downstream objective, tailored loss functions are employed to ensure optimal task-specific adaptation. The implementation details are provided in Section 4.2.
Figure 2. The pre-trained encoder is connected to a linear prediction head to perform downstream tasks.
3.3. EEG Data Generation by Diffusion Model
As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging. To address this, we utilize PCA to project raw data into a high-SNR subspace, establishing a refined latent manifold as the denoising objective. Similar to the LDMs (Rombach et al., 2022), we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals. In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process. After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
3.3.1. Diffusion Formulation
The input to our proposed latent diffusion model, which is conditioned on channel augmentation, is the latent representation $z=Px$ obtained through PCA pre-processing. Here, we use the $\omega$-length window to segment each $x$ before applying PCA to obtain the corresponding latent representation. The PCA basis vectors are computed via eigen-decomposition without the need for gradient-based training. Our diffusion model operates within this latent space. The noisy input $z_{t}$ at a given timestep is transformed into patches, which are then linearly embedded into a sequence of tokens with dimension $d$ (implemented by the patchify layer). Here, we apply a sine-cosine version of the positional embeddings to all input tokens. Then, the input tokens are processed by several DiT blocks. By applying the reparameterization trick, we can do the sampling as follows:
(1) $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$
where $\epsilon_{t}\sim{\mathcal{N}(0,\mathbf{I})}$. The diffusion model is trained to learn the reverse denoising process:
(2) $p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(\mu_{\theta}(z_{t}),\Sigma_{\theta}(z_{t})),$
where the network is used to predict the statistics of $p_{\theta}$. By reparameterizing $\mu_{\theta}$ as a noise prediction network (i.e., denoiser) $\epsilon_{\theta}$, the model can be trained using the mean squared error between the predicted noise $\epsilon_{\theta}(z_{t})$ and the ground truth sampled Gaussian noise $\epsilon_{t}$:
(3) $\mathcal{L}_{simple}(\theta)={\left\|{\epsilon_{\theta}(z_{t})-\epsilon_{t}}\right\|}_{2}^{2}.$
To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al. (Nichol and Dhariwal, 2021): train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$. After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction. Both outputs have the same data shape as the input. Once $p_{\theta}$ is trained, new latent representations can be sampled by initializing $z_{t_{max}}\sim{\mathcal{N}(0,\mathbf{I})}$ and sampling $z_{t-1}\sim{p_{\theta}(z_{t-1}|z_{t})}$ via the reparameterization trick. Finally, we obtain the denoised latent representation $\widetilde{z}$, which is then projected back to the original signal space using inverse PCA to reconstruct the original EEG signals:
(4) $\widetilde{x}=P^{-1}\widetilde{z}=P^{T}\widetilde{z}$.
3.3.2. Conditional Guidance
We incorporate the encoded EEG representation as conditional input for the diffusion model. It guides the generative process and is further leveraged for downstream tasks. To this end, we broadly consider EEG signals in this context as any collection of EEG recordings that exhibit meaningful interrelationships (e.g., semantic correlations), particularly those derived through various augmentations of source EEG signals — an approach well-established in contrastive learning frameworks (Chen et al., 2020). Specifically, each augmented EEG signal is encoded using a shared EEG encoder, and the resulting EEG representations $e^{1:m}$ are aggregated into a single vector $e$ by taking their mean, enabling the model to incorporate richer conditional information. When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$. Under this setting, classifier-free guidance (Ho and Salimans, 2022) can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized. According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$. By interpreting the output of the diffusion model as a score function, the sampling process can be guided by:
(5) $\widetilde{\epsilon}_{\theta}(z_{t},e)\propto{\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\epsilon_{\theta}(z_{t},e)-\epsilon_{\theta}(z_{t},\emptyset))}},$
$\widetilde{\epsilon}_{\theta}(z_{t},e)=\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\nabla_{z}\log p(z|e))},$
where $s>1$ denotes the scale of the guidance. The diffusion model with $e=\emptyset$ is trained by randomly dropping $e$ during training and replacing it with a learned null embedding $\emptyset$. The classifier-free guidance sampling method can significantly improve the quality of generated samples by effectively utilizing the conditional information we introduced.
3.3.3. Conditional Input Setup
The treatment of conditional inputs and the initialization process significantly influence the generation quality in diffusion models. We adopt the adaLN-Zero block from DiT, which has shown superior performance, to handle the conditional inputs (the EEG representations $e$ and the noise timestep $t$). Specifically, we regress the scale and shift parameters $\gamma$ and $\beta$ in the Transformer blocks based on the sum of the $e$ and $t$. Additionally, we regress the dimension-wise scaling parameters $\rho$, which are applied just before each residual connection in the DiT blocks. The MLP for all $\rho$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function. In addition to layer normalization, we incorporate conditional embeddings into residual connections, which enhance the vanilla adaLN and improve the quality of the generated EEG signals.
4. Experiment
4.1. Datasets
Pre-training datasets. We utilize the Temple University Hospital EEG Corpus (TUEG) (Obeid and Picone, 2016) to pre-train the proposed EEGDM. TUEG is a large-scale publicly available clinical EEG dataset containing clinical-grade recordings collected in real-world hospital environments. It encompasses over 40 different channel configurations and recordings of varying temporal durations, with the majority acquired at a sampling rate of 256 Hz. TUEG exhibits substantial heterogeneity in patient demographics, channel configurations, and recording equipment, making it particularly suitable for pre-training our model.
Downstream tasks and Datasets. To comprehensively evaluate our method, we consider multiple downstream tasks, including event type classification (TUEV dataset (Harati et al., 2015)), abnormal detection (TUAB dataset (Lopez et al., 2015)), motor imagery classification (SHU-MI dataset (Ma et al., 2022)), cognitive load assessment (CL-Drive dataset (Angkan et al., 2024)), emotion recognition (Seed-VII dataset (Jiang et al., 2024b)), and emotion distribution prediction (DMER dataset (Yang et al., 2024b)). Detailed descriptions of these datasets are provided in the supplementary material.
Notably, we primarily select downstream datasets from the past five years, representing a strategic shift from the benchmarks used in prior works such as CBraMod (Wang et al., 2025) and EEGPT (Wang et al., 2024). This choice is motivated by the need to evaluate the model against the fidelity and complexity of contemporary neuroscience standards. Older datasets often lack the high spatial density and signal quality required to stress-test a foundation model, whereas recent paradigms—ranging from naturalistic stimuli to multimodal retrieval—offer a more effective assessment of the model’s capacity for generalizable, high-level cognitive decoding. Nevertheless, to ensure fair comparison with established baselines, we make exceptions for the TUAB and TUEV datasets. Although these two widely adopted benchmarks fall outside the five-year window, they are retained as standardized anchors, allowing us to validate our model’s performance on classical clinical tasks alongside its capabilities on modern, complex scenarios.
4.2. Experimental Setup
Data Pre-processing. For the pre-training dataset TUEG, following prior work (Wang et al., 2025), we select 19 common EEG channels. All EEG signals are resampled to 200 Hz and segmented into non-overlapping 30-second samples (19 channels $\times$ 6,000 time points). Each EEG patch is set to a temporal length of 1 second, resulting in 570 patches per sample. For all downstream datasets, we maintain consistency with our pre-training setup by resampling EEG signals to 200 Hz and setting each EEG patch duration to 1 second.
Table 1. Performance comparison on TUEV. The best result is highlighted in bold, and the second-best result is underlined.
Methods Params Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
[…]
Table 2. Performance comparison on TUAB. The best result is highlighted in bold, and the second-best result is underlined. NA indicates that the result is not available in the original literature (Wang et al., 2024).
Methods Params Balanced Accuracy ↑ AUROC ↑ AUC-PR ↑
[…]
Baselines and Implementation. We compare our EEGDM against state-of-the-art self-supervised EEG representation learning frameworks, including EEGPT (Wang et al., 2024), LaBraM-Base (Jiang et al., 2024c) (the only variant with available pre-trained weights), CBraMod (Wang et al., 2025), and Puah-EEGDM-SSM (Puah et al., 2025). All baselines are fine-tuned using their publicly available code and pre-trained weights, except in cases where results were directly reported in the original papers. For our model, we adopt the DiT architecture for the EEG generator (diffusion module) and a ViT backbone for the encoder. The model is pre-trained on a cluster of four NVIDIA A100 GPUs with a batch size of 32. The entire pre-training phase encompasses 329,400 steps, totaling approximately 112.96 wall-clock hours.
Evaluation Protocols. We evaluate our method under two settings: fine-tuning and linear probing (with a frozen encoder), to comprehensively assess the quality of the learned representations. Extensive experiments are conducted to evaluate the proposed model, demonstrating its strong representation capability across diverse downstream datasets.
Detailed hyperparameter configurations, data splitting strategies, and implementation specifics for both pre-training and downstream evaluation protocols are provided in the supplementary material.
Table 3. Performance comparison on DMER. The best result is highlighted in bold, and the second-best result is underlined.
Methods Chebyshev ↓ Clark ↓ Canberra ↓ KL ↓ Cosine ↑ Intersection ↑
[…]
Table 4. Performance comparison on SHU-MI, CL-Drive and Seed-VII datasets. The best result is highlighted in bold, and the second-best result is underlined. Results for Puah-EEGDM-SSM on SHU-MI are omitted due to performance near chance levels.
Datasets Methods Balanced Accuracy ↑ Weighted F1 / AUROC ↑ Cohen’s Kappa / AUC-PR ↑
[…]
4.3. Results and Analysis
In this section, we demonstrate that EEGDM achieves competitive performance across diverse downstream tasks, underscoring the effectiveness of diffusion-based representation learning for EEG analysis. All experiments are conducted with multiple random seeds, and we report the mean and standard deviation of evaluation metrics. A comprehensive evaluation reveals that among all compared methods, our EEGDM is the only model that attains a top-two position on all six downstream datasets (except for EEGDM (T-Pre) on TUEV) with respect to the primary evaluation metrics (balanced accuracy and KL divergence).
As shown in Table 1, EEGDM achieves a balanced accuracy of 68.19% on the TUEV dataset when pre-trained on the large-scale TUEG corpus, outperforming several state-of-the-art baselines. Notably, while the general-purpose version of EEGDM trails the results reported by Puah-EEGDM-SSM (Puah et al., 2025) (75.57%) on this specific benchmark, it is essential to consider the difference in pre-training data distributions. Unlike our model, which targets universal feature extraction across diverse domains (as evidenced by its superior results across other downstream datasets), Puah-EEGDM-SSM was pre-trained and evaluated exclusively on the TUEV dataset, benefiting from a high degree of domain alignment. To ensure a rigorous and fair architectural comparison, we introduced a task-specific variant of our model, denoted as EEGDM (T-Pre), which follows the same protocol as Puah-EEGDM-SSM (Puah et al., 2025) by utilizing TUEV for both pre-training and downstream evaluation. Under this aligned setting, EEGDM (T-Pre) achieves a state-of-the-art accuracy of 79.71%. This result demonstrates the superior representational capacity of our Latent Diffusion Transformer (DiT) architecture operating on a refined manifold. It further indicates that the performance gap in the general setting is a byproduct of domain shift in pre-training data rather than architectural limitations. The discrepancy between balanced accuracy and F1-score suggests that EEGDM’s latent space prioritizes holistic representation over majority-class optimization, leading to better pattern diversity at the cost of higher sensitivity to class boundary noise.
Table 2 shows that EEGDM achieves a balanced accuracy of 81.80% and an AUC-PR of 90.85% on the TUAB dataset. Securing the second-best performance on those metrics, EEGDM demonstrates competitive performance with a narrow margin. The consistent performance across both balanced accuracy and AUC-PR metrics suggests that EEGDM learns robust representations that generalize well to binary abnormality detection tasks, maintaining high discriminative power even in challenging clinical scenarios.
Table 3 presents the results for the DMER dataset, which uses distributional metrics. EEGDM achieves the best performance across various divergence and similarity metrics, with the exception of the Clark metric, where it ranks second. The strong performance on DMER, a task requiring the prediction of an emotion distribution rather than a discrete class, strongly suggests that diffusion-based pre-training learns rich, fine-grained features that preserve the complex, subtle, and continuous emotional states embedded in EEG signals.
Table 4 further validates the generalizability and robustness of EEGDM across diverse BCI tasks. Specifically, on the SHU-MI dataset, EEGDM ranks second in balanced accuracy (63.24%) and AUC-PR (70.70%), while achieving the best performance in AUROC (69.96%), demonstrating competitive performance compared with CBraMod. Notably, the results for Puah-EEGDM-SSM (Puah et al., 2025) on this specific benchmark are omitted from the table, as its performance fluctuated near chance levels (e.g., accuracy $\approx 50\%$). This discrepancy suggests that while domain-specific pre-training (as used in (Puah et al., 2025)) may excel in aligned tasks, our Latent Diffusion Transformer (DiT) architecture, pre-trained on a broader corpus, exhibits superior cross-task adaptability, particularly in capturing the subtle spatial-temporal patterns required for motor imagery. On the CL-Drive dataset, EEGDM achieves the best balanced accuracy and the best Cohen’s Kappa, demonstrating superior cognitive load feature extraction capabilities. On the Seed-VII dataset, EEGDM continues its strong performance with the best balanced accuracy and ranks second in weighted F1 and Cohen’s Kappa, further validating its effectiveness in emotion-related tasks.
Table 5. Performance comparison of two data augmentation strategies on the CL-Drive dataset.
Strategy Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
[…]
Beyond representation learning, a distinctive advantage of our proposed EEGDM is its generative capability to synthesize high-fidelity EEG data, an ability inherently absent in mask-reconstruction frameworks such as MAE. This ability is particularly effective in mitigating the pervasive issues of data sparsity and class imbalance in neural decoding. We evaluate this capability on the CL-Drive dataset—a relatively small-scale benchmark characterized by significant class skewness—using two augmentation strategies: isoportion expansion and balanced augmentation. As summarized in Table 5, our generative augmentation significantly improves downstream performance. By populating under-represented manifolds with synthetic yet semantically consistent samples, EEGDM effectively rectifies biased decision boundaries caused by data scarcity. These results underscore that our diffusion-based pre-training not only distills robust features but also serves as a powerful plug-and-play data engine, enhancing the model’s adaptability to real-world, long-tailed EEG distributions.
Comprehensive qualitative evaluations, including t-SNE visualization and EEG signal reconstruction, are provided in the supplementary material to further validate the representation quality and generative fidelity of EEGDM.
4.4. Ablation Study
To verify the individual and collective contributions of PCA and channel augmentation, we evaluate four variants of EEGDM, as summarized in Table 6. Our analysis reveals a compelling synergistic relationship between these two components that extends beyond simple additive effects. Interestingly, the results indicate that employing either PCA or channel augmentation in isolation provides marginal improvements or even leads to slight performance degradation. While PCA projects signals into a latent space with improved signal-to-noise ratio (SNR), the reduction in dimensionality may inadvertently discard fine-grained temporal nuances. Without sufficient input diversity, the encoder tends to underfit the complex underlying brain states. Conversely, although channel augmentation enriches the training perspectives, it significantly increases the optimization burden when the model is tasked with reconstructing raw, noisy signals.
The state-of-the-art performance achieved by Variant D underscores a powerful complementary effect. PCA simplifies the latent manifold, making the model more sensitive to the structural information provided by the augmented views. Meanwhile, the diverse augmented inputs prevent the model from losing critical information—effectively compensating for the potential information loss incurred by dimensionality reduction. To further assess the generalizability of this mechanism, we conducted a low-resource fine-tuning experiment on the TUAB dataset using only 10% of the training data (last column); we also use this setting in Fig. 3. The results consistently show that the integration of PCA and EEG augmentation remains vital even in small-scale data settings.
Table 6. The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
Variants PCA EEG Augmentation TUEV TUAB (10%)
[…]
(a) TUEV
(b) TUAB
Figure 3. Performance comparison with different numbers of PCA components on TUEV and TUAB datasets.
We also investigate the impact of the number of PCA components on downstream task performance. Fig. 3 presents results on TUEV and TUAB with varying numbers of PCA components, revealing a non-monotonic relationship between the number of PCA components and downstream task performance. Specifically, an insufficient number of components may result in substantial information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance. To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.
In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG generation, as validated by its effectiveness in augmenting an imbalanced and small-scale dataset; (2) Effective representation learning, as demonstrated by competitive linear probing results and substantial performance gains through fine-tuning; and (3) Competitive performance across diverse downstream tasks, as evidenced by consistent improvements in balanced accuracy and robustness over state-of-the-art baselines.
5. Conclusion
We present EEGDM, a novel latent diffusion framework for EEG representation learning, through explicit conditioning on EEG channel augmentation. Our approach adopts the generative task as the self-supervised EEG representation learning objective. By (1) integrating a denoiser with the EEG encoder and (2) leveraging PCA-based projection and EEG channel augmentation, EEGDM performs the generation task in the latent space using rich conditional information while simultaneously learning powerful EEG semantic representations that can be applied to diverse downstream tasks. Comprehensive experimental results demonstrate the model’s effectiveness across multiple downstream tasks.
Supplementary Material
Appendix A Dataset Details for Downstream Tasks
This section provides a detailed description of the downstream datasets employed to evaluate the performance of our proposed model.
(1) TUEV (event type classification) (Harati et al., 2015): An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). The EEG signals are recorded with 23 channels at a sampling rate of 256 Hz.
(2) TUAB (abnormal detection) (Lopez et al., 2015): A corpus of EEG recordings annotated as either normal or abnormal. Similar to TUEV, the EEG signals in TUAB are recorded with 23 channels at a sampling rate of 256 Hz.
(3) SHU-MI (motor imagery classification) (Ma et al., 2022): An EEG motor imagery dataset designed for decoding imagined movements. It features 32 EEG channels sampled at 250 Hz, including two types of motor imagery tasks: left hand and right hand.
(4) CL-Drive (cognitive load assessment) (Angkan et al., 2024): A multimodal dataset designed for cognitive load assessment, which includes EEG signals. The EEG data are recorded using four channels at a sampling frequency of 256 Hz. This study explores the ternary classification of cognitive load into high, medium, and low levels.
(5) Seed-VII (emotion recognition) (Jiang et al., 2024b): A large-scale multimodal emotion dataset. Seed-VII includes 62-channel EEG and eye-movement data from 20 healthy subjects across seven emotion categories (disgust, fear, sad, neutral, happy, anger, surprise). The preprocessed EEG signals were downsampled to 200 Hz.
(6) DMER (emotion distribution prediction) (Yang et al., 2024b): A multimodal emotion recognition corpus that includes EEG signals along with other modalities (GSR, PPG, and frontal facial videos). In each trial, participants watched video clips and subsequently completed a self-assessment using a ten-item version of the Positive Affect and Negative Affect Schedule (PANAS) (Mackinnon et al., 1999), reporting the intensity of ten emotional adjectives. Each emotion rating ranges from 1 (not at all) to 5 (extremely), and the resulting scores are transformed into a 10-dimensional emotion distribution. The EEG recordings comprise 18 channels sampled at 300 Hz.
Appendix B Detailed Experimental Settings for Downstream Datasets
For the downstream datasets TUEV and TUAB, consistent with previous studies (Wang et al., 2024; Jiang et al., 2024c; Wang et al., 2025), we utilize 16 common bipolar montage channels from the international 10-20 system, with sample durations of 5 seconds and 10 seconds, respectively. We adopt the train-test splits provided by these datasets. For the CL-Drive dataset, we use the channel configuration and sample duration (10 seconds) provided by the dataset and apply 10-fold cross-validation. For the SHU-MI, DMER, and SEED-VII datasets, we use the channel configurations provided by each dataset and segment the signals into 4-second samples. A subject-independent split strategy is adopted for these datasets.
Appendix C Implementation Details
Table 7. Hyperparameters for EEGDM pre-training.
Hyperparameter Value
Optimization Batch size 32
Peak learning rate 5e-4
Minimal learning rate 5e-5
Learning rate scheduler OneCycleLR
Weight decay 0.05
EMA decay rate 0.9999
Gradient clipping norm 1
Optimizer AdamW
$\beta$} (0.9, 0.999)
Diffusion Module Layers N 8
Hidden size 512
Heads 8
Normalization type LayerNorm
EEG Encoder Layers 8
Hidden size 512
MLP size 2048
Heads 8
Sampling Classifier-free guidance 9
Diffusion training steps 1000
Pre-training: The pre-training of EEGDM is conducted on the large-scale TUEG dataset to learn robust, general-purpose neural representations. We utilize the AdamW optimizer with a peak learning rate of 5e-4, which is decayed to a minimal value of 5e-5 following a OneCycleLR scheduling policy. We apply an Exponential Moving Average (EMA) on the model weights with a decay rate of 0.9999. The architecture consists of an 8-layer EEG Encoder and an 8-layer Diffusion Module, both featuring a hidden size of 512 and 8 attention heads. Detailed hyperparameter configurations for the optimization process, model architecture, and diffusion sampling are summarized in Table 7.
Downstream tasks: We evaluate the pre-trained EEGDM across diverse downstream scenarios using two protocols: fine-tuning and linear probing. In the fine-tuning setting, both the pre-trained encoder and the task-specific prediction head are jointly optimized. For linear probing, the EEG encoder remains strictly frozen, and only the prediction head is trained to assess the quality of the learned representations. To accommodate the varying scales and complexities of different downstream datasets, we employ task-specific hyperparameter configurations, as detailed in Table 8. All downstream experiments are conducted using the AdamW optimizer to ensure training stability. Detailed hyperparameters for both protocols, including batch sizes, learning rates, and weight decay, are summarized in Table 8.
Table 8. Hyperparameters for downstream fine-tuning and linear probing. ($\star$) Depends on the dataset.
Hyperparameter Fine-tuning Linear Probing
Encoder status Unfrozen Frozen
Learning rate 5e-4 2e-4, 5e-4, 2e-3
Total epochs 30, 150 30, 50, 100
Weight decay 0.05 0.05, 0.02, 0.01
Batch size 64 32, 64
Optimizer AdamW AdamW
Learning rate scheduler OneCycleLR OneCycleLR, ReduceLROnPlateau, CosineAnnealingLR
$\beta$} (0.9, 0.999) (0.9, 0.999)
Warmup epochs 6, 30 3, 10
Appendix D Qualitative Analysis
To evaluate the quality of the learned representations, we visualize the EEG data embeddings using t-SNE (Van der Maaten and Hinton, 2008) as presented in Fig. 4. We randomly sample EEG signals from the test sets of two downstream tasks in a class-balanced manner and extract their representations using our pre-trained encoder. The t-SNE visualizations exhibit well-separated clusters across all datasets, indicating that our EEGDM learns discriminative representations capable of capturing task-relevant patterns. This clear clustering structure further demonstrates the EEGDM’s ability to learn generic yet informative EEG representations that can be successfully adapted to diverse downstream tasks.
To further validate the generative quality, we provide visualizations of reconstructed signals and their Power Spectral Density (PSD). As shown in Fig. 5, the reconstructed EEG signals on the pre-training dataset, as well as their frequency-domain representations, closely match the ground truth. While minor discrepancies remain at the fine-grained level, the global temporal structure and frequency characteristics are well preserved. These results indicate that the proposed model can generate realistic EEG signals and effectively capture high-level semantic information from EEG data.
(a) TUEV
(b) TUAB
Figure 4. The t-SNE visualizations. Clear inter-class separation and intra-class compactness indicate that the learned representations are discriminative.
Figure 5. Visualization of reconstructed EEG signals and their PSD.